1. A tank in the shape of an inverted pyramid has a square base of 4 m on a side and an altitude of 6 m. How many liters of water will it hold when full?
1
Expert's answer
2020-12-07T18:39:12-0500

The volume of the pyramid is V=1/3Bh

B is the area of the base, h is the altitude and h=6 m. B=a2, a=4, B=16m2.

Volume V=1/3*16*6=32 m3